OpenStudy (anonymous):

Can someone please check my work? I'll post my work and answer in the comments. I'm not sure if I completed it properly.

OpenStudy (anonymous):

Find the solutions of [2x ^{2}-16x+32=0\]

OpenStudy (anonymous):

My work: \[x= 16\pm \frac{\sqrt{16^2}-4(2)(32) }{ 2(2) }\] \[x= 16\pm \frac{ 256-4(64) }{ 4 }\] \[x= 16\pm \frac{\sqrt{256 -256}}{ 4 }\] \[x= 16\pm \sqrt{0}\] \[x= \frac{ 16+0 }{ 4 } , \frac{ 16-0 }{ 4 } = (0,0)\]

OpenStudy (anonymous):

If x = 0 then that should mean there are no solutions, correct?

OpenStudy (anonymous):

Because when I check my work I get: \[2(0)^2-16(0)+32\] \[0-0+32=0\] \[32 \neq 0\]

OpenStudy (johnweldon1993):

Where did the 0 come from? \[\large x = \frac{16 \pm 0}{4} \cancel{=} 0\]

OpenStudy (anonymous):

When I took x=16±√0 and separated the addition and subtraction symbols so I would have two equations \[x= \frac{ 16+0 }{ 4 } , \frac{ 16-0 }{ 4 }\]

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

I see what I did wrong, it would \[\frac{ 16 }{ 4 }\] regardless of addition or subtraction

OpenStudy (johnweldon1993):

There ya go!

OpenStudy (anonymous):

Thank you so much! I can't believe I missed something so simple.

OpenStudy (johnweldon1993):

Lol it happens dont worry about it!